Suppose you are investigating the reaction: M(s) + 2 HCl(aq) → MCl2(aq) + H2(g). You weigh out a 0.202 gram piece of metal and combine it with 76.3 mL of 1.00 M HCl in a coffee-cup calorimeter. If the molar mass of the metal is 48.37 g/mol, and you measure that the reaction absorbed 121 J of heat, what is the enthalpy of this reaction in kJ per mole of limiting reactant?

Enter your answer numerically to three significant figures in units of kJ/mol.
Answer #1

no of moles of M = W/G.M.M

                            = 0.202/48.37

                            = 0.004176moles

no of moles of HCl   = molarity * volume in L

                                = 1*0.0763   = 0.0763 moles

M(s) + 2 HCl(aq) → MCl2(aq) + H2(g)

1 mole of M react with 2 moles of HCl

0.004176 moles of M react with = 2*0.004176/1   = 0.008352moles of HCl is required

HCl is limiting reactant

\DeltaH   = 121J/0.008352 Mole

         = 14488J/mole

         = 14.488KJ/mole

        round off

         = 14.5KJ/mole >>>>answer